OpenStudy (anonymous):

Calculate the integral of cos^4(3x)dx..Need some help

OpenStudy (anonymous):

I got you :)

OpenStudy (amistre64):

gonna need a 3 in there somplace ;)

OpenStudy (amistre64):

and a -sin ... right?

OpenStudy (anonymous):

cos^2 x - cos^2 x sin^2 x

OpenStudy (anonymous):

in place of x put 3x

OpenStudy (anonymous):

right ... whats the formula for that..

OpenStudy (anonymous):

convert cos^2 3x into (1+ cos6x)/2 and cos^2 3x sin^2 3x into 0.5sin^2 (6x)

OpenStudy (anonymous):

you just do the same thing thing as a normal cos^2(x) just replace x with 3x?

OpenStudy (anonymous):

Rewrite it as: \[\int\limits (\cos^2(3x))^2 dx\] You want to use your half angle formula: \[\cos^2(x)=\frac{1}{2}+\frac{1}{2}\cos(2x)\] Then: \[\int\limits (\frac{1}{2}+\frac{1}{2}\cos(6x))^2dx\] Square it out, then use the identity again. Integrate term by term. Or use a power reduction formula >.>

OpenStudy (anonymous):

1/2 (sin^2 (6x)) = 1/4(1- cos12x)

OpenStudy (anonymous):

u can convert like ive done and then integrate

OpenStudy (anonymous):

whats teh half angle formula ?

OpenStudy (anonymous):

whats teh half angle formula ?

OpenStudy (anonymous):

What I used^^

OpenStudy (anonymous):

More generally: \[\cos^2(ax)=\frac{1}{2}+\frac{1}{2}\cos(2ax)\]

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

Did you follow what I did above? I can finish working it out if need be.

OpenStudy (anonymous):

why did you write as (cos^2(3x))^2 dx?

OpenStudy (anonymous):

So that you could see the formula I gave 3 posts up is applicable (sometimes its hard to see with higher powers) but notice that (cos^2(3x))^2=cos^4(3x). Does that make sense?

OpenStudy (anonymous):

yes, you converted it just so you can use the double angle formula?

OpenStudy (anonymous):

Yes :) So going from what I have above: \[\int\limits (\frac{1}{2}+\frac{1}{2}\cos(6x))^2dx=\int\limits (\frac{1}{4}+\frac{1}{2}\cos(6x)+\frac{1}{4}\cos^2(6x))dx\] Do you follow that distribution?

OpenStudy (anonymous):

Refresh if the latex looks bad*

OpenStudy (anonymous):

Yes i got that far

OpenStudy (anonymous):

Okay then use the identity again on the cos^2(6x) giving: \[\int\limits(\frac{1}{4}+\frac{1}{2}\cos(6x)+\frac{1}{4}(\frac{1}{2}+\frac{1}{2}\cos(12x)))dx\] Distribute the 1/4. From there do you know how to proceed?

OpenStudy (anonymous):

so its the integration of (3/8) +(1/2)cos(6x)+(1/8)cos(12x))dx

OpenStudy (anonymous):

Exactly :) What did you get when you did that?

OpenStudy (anonymous):

i would split the integration right?

OpenStudy (anonymous):

You could. Giving: \[\int\limits \frac{3}{8} dx+\frac{1}{2} \int\limits \cos(6x)dx+\frac{1}{8} \int\limits \cos(12x)dx\] If that helps you.

OpenStudy (anonymous):

i got (1/8)sin(6x)+ (1/96)sin)(12x)

OpenStudy (anonymous):

That's right, almost. Don't forget you have the integral of 3/8 which is (3/8)x. So you need a +(3/8)x too :P

OpenStudy (anonymous):

And a +c to be "technically" right xD

OpenStudy (anonymous):

oh ya

OpenStudy (anonymous):

But good job :) Any time you have powers of cosine and sine and they are both EVEN. You will use that identity to simplify it. The one for sine is the same just off by a negative: \[\sin^2(ax)=\frac{1}{2}-\frac{1}{2}\sin(2ax)\]. When you start mixing even and odd powers you'll need: \[\sin^2(x)+\cos^2(x)=1\] Its dependent on a problem-by-problem basis. Let me know if you get stuck on another one :P

OpenStudy (anonymous):

Errrr...that should read sin^2(ax)=1/2-(1/2)COS(2ax). Mistype on my part.

OpenStudy (anonymous):

thanks so much .. i definitely will

OpenStudy (anonymous):

No problem :) I'll be around for a while

OpenStudy (anonymous):

its telling me its incorrect ..

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+cos^4(3x) It's correct (just distribute the 1/96)

OpenStudy (anonymous):

I don't know if it won't take that, me, you, and wolfram got the same thing...

OpenStudy (anonymous):

this site is awesome..

OpenStudy (anonymous):

It will do any integration for you. Then just click: see steps. However, if you do definite integrals, it won't show you the step by step. So just do indefinite then check your answer, then do the numbers.

OpenStudy (anonymous):

Also does any math anything. Or type in something like GDP ireland 2005. It'll give you that. Type in radius of the moon or ANYTHING.

OpenStudy (anonymous):

i dont get how they got a final anwer from (3/8)x-(3/32)sin(4x)-(1/8)sin^3(2x)cos(2x)+C ..of (1/64)(24x-8sin(4x)+sin(4x)) +C